OpenStudy (princess_shadow_sob):

Circle A has a diameter of 9 inches, a circumference of 28.26 inches, and an area of 63.585 square inches. The diameter of circle B is 5 inches, the circumference is 15.70 inches, and the area is 19.625 square inches.. Part A: Using the formula for circumference, solve for the value of pi for each circle. (4 points) Part B: Use the formula for area and solve for the value of pi for each circle. (4 points) Part C: What observation can you make about the value of pi for circles A and B? (2 points)

OpenStudy (princess_shadow_sob):

@Eboogie4 @Error1603

OpenStudy (princess_shadow_sob):

@suppa_lova

OpenStudy (eboogie4):

@-SpencerBraz- @Qwertty123

OpenStudy (anonymous):

ok so what are the formulas that you need to use please write them out bc it has been a while

OpenStudy (error1603):

1st wat u want to do is Divide 28.26 with 9 and u will get PI which is 3.14

OpenStudy (princess_shadow_sob):

C=2πr

OpenStudy (anonymous):

3.14159265 to be exact

OpenStudy (princess_shadow_sob):

no its 3.14 suppa_lova so @Error1603 whats next

OpenStudy (error1603):

2nd you want to Divide 15.70 by 5 and also you will get PI which is 3.14 again

OpenStudy (error1603):

3rd divide 63.585 by 20.25 and again you will get PI which is 3.14

OpenStudy (error1603):

Got it?

OpenStudy (princess_shadow_sob):

yep

OpenStudy (princess_shadow_sob):

now what

OpenStudy (error1603):

Oh, we just solved it. You just simply put it all together, like this: A = 3.14 B = 3.14 C =The person working these calculations for circles A and B is using the same consistent value for pi or 3.14

OpenStudy (princess_shadow_sob):

so whats the circumference

OpenStudy (princess_shadow_sob):

oh k

OpenStudy (princess_shadow_sob):

stay i have another one

OpenStudy (error1603):

Okay

OpenStudy (princess_shadow_sob):

hold on i coping the stuff just had to take care of my bird

OpenStudy (error1603):

Alright

OpenStudy (error1603):

Sorry got to go

OpenStudy (princess_shadow_sob):

@xMissAlyCatx

OpenStudy (princess_shadow_sob):

A candy bar box is in the shape of a triangular prism. The volume of the box is 3,240 cubic centimeters. A triangular prism is shown with base of triangle labeled 24 cm, sides of the triangle labeled 15 cm, and length of the box equal to 30 cm. Part A: What is the height of the box? Show your work. (5 points) Part B: What is the approximate amount of cardboard used to make this candy box? Explain how you got your answer. (5 points)

OpenStudy (princess_shadow_sob):

can u help me @xMissAlyCatx

OpenStudy (xmissalycatx):

Yes. Hello there :)

OpenStudy (princess_shadow_sob):

hello

OpenStudy (xmissalycatx):

Okay, so do you know the formula for the volume of a triangular prism?

OpenStudy (princess_shadow_sob):

google says V=1 4h﹣a4+2(ab)2+2(ac)2﹣b4+2(bc)2﹣c4

OpenStudy (princess_shadow_sob):

V≈4949.18

OpenStudy (princess_shadow_sob):

thats what google said when i put the # in

OpenStudy (xmissalycatx):

The volume for a triangular prism is: \[V=\frac{ 1 }{ 2 }*b*h*l\]

OpenStudy (princess_shadow_sob):

when i wnet to google it said differently but ok

OpenStudy (princess_shadow_sob):

so whats the answer i always get this stuff wrong

OpenStudy (xmissalycatx):

That's the formula, I promise. Okay anyways.. What you want to do is plug in the numbers that it gave to you on the picture and in the beginning into to formula..

OpenStudy (xmissalycatx):

Do that for me please.

OpenStudy (princess_shadow_sob):

what i dont under stand

OpenStudy (xmissalycatx):

Okay, the numbers that you were given on your picture.. plug them into that formula I just gave you..

OpenStudy (princess_shadow_sob):

i dont know how to plug them in the formula

OpenStudy (xmissalycatx):

\[3240=\frac{ 1 }{ 2}*30*24*h\]

OpenStudy (xmissalycatx):

So multiply the 1/2 the 30 and the 24..

OpenStudy (princess_shadow_sob):

360 I DONT KNOW i dont know how to but a 1/2 * 30* 24 in the calculator

OpenStudy (xmissalycatx):

Just do 30 x 24 x .5

OpenStudy (xmissalycatx):

.5 is the same as 1/2

OpenStudy (princess_shadow_sob):

so 360 like i said

OpenStudy (xmissalycatx):

Yes.

OpenStudy (princess_shadow_sob):

so now what

OpenStudy (xmissalycatx):

Okay so now you've got: 3240 = 360 h So you need to divide 360 from both sides so that "h" is by itself.

OpenStudy (princess_shadow_sob):

so .5 bc if u divide 360/30/24=.5

OpenStudy (xmissalycatx):

.5 is what 1/2 is.. that's where you get .5 from.

OpenStudy (xmissalycatx):

What is 3240/360?

OpenStudy (princess_shadow_sob):

9

OpenStudy (xmissalycatx):

That's your answer. h=9

OpenStudy (princess_shadow_sob):

so what B

OpenStudy (xmissalycatx):

Are you familiar with finding surface area of each part of the prism?

OpenStudy (princess_shadow_sob):

i think so @xMissAlyCatx

OpenStudy (xmissalycatx):

Okay. So find the surface area of one of the bases, and one of the faces..

OpenStudy (princess_shadow_sob):

A≈1415.89 a Base side 24 b Base side 30 c Base side 30 h Height 9

OpenStudy (princess_shadow_sob):

i dont know help me plz @xMissAlyCatx

OpenStudy (xmissalycatx):

Okay, let's just look at the triangle base first.. We have the Area for a triangle which is A = 1/2 times b times (h)

OpenStudy (xmissalycatx):

So A = 1/2 (24) (9)

OpenStudy (princess_shadow_sob):

so .5*9*24=108

OpenStudy (xmissalycatx):

Yes.

OpenStudy (xmissalycatx):

Okay, now multiply that by 2 because you have 2 of the bases in total.

OpenStudy (princess_shadow_sob):

so the answer for question B is108

OpenStudy (xmissalycatx):

No. We aren't done yet.

OpenStudy (xmissalycatx):

You need the surface area of the 3 faces as well.

OpenStudy (xmissalycatx):

30 (15) + __?__

OpenStudy (xmissalycatx):

=** not + sorry

OpenStudy (princess_shadow_sob):

108*2=216

OpenStudy (xmissalycatx):

Yes for that part. Now do the 30 (15) = ?

OpenStudy (princess_shadow_sob):

30*15=450

OpenStudy (xmissalycatx):

Okay good. Now multiply that by 3.

OpenStudy (princess_shadow_sob):

450*3=1350

OpenStudy (xmissalycatx):

Okay, now add 1350 and 216..

OpenStudy (princess_shadow_sob):

so 1566

OpenStudy (xmissalycatx):

Yes. It would be 1566 cm squared.